OpenStudy (anonymous):

If interest paid on an investment is $25.21 for 50 days at an interest rate of 8.25%, what was the original amount invested if the exact interest method is used? (Round to the nearest cent).

OpenStudy (ash2326):

Interest Paid= $25.21 Time= 50 days Rate= 8.25 % In exact Interest method 1 year= 365 days so time= 50/365 so we need to find the original amount invested , Let it be P \[Interest = \frac{P \times R \times T}{100}\] \[25.21=\frac{P\times 8.25 \times \frac{50}{365}}{100}\] So \[P= \frac{25.21 \times 100 \times 365}{8.25 \times 50}\] Now you can find P, can't you?

OpenStudy (ash2326):

Hi @acooper8525 Welcome to Open Study:D

OpenStudy (anonymous):

hi and thanks

OpenStudy (ash2326):

Did you get it?

OpenStudy (anonymous):

working on it now

OpenStudy (ash2326):

Great:)

OpenStudy (anonymous):

P=2,230.70?

OpenStudy (anonymous):

?